# Chinese Academy of Sciences, Beijing 100190, P. R. China

### v1

@@ 9 Constraints and implications from PBH formation
Most of the domain wall energy $\rho_{w}$ is released at the annihilation temperature $T_{\rm ann}$, beyond which $\rho_{w}$ scales as $T^{\alpha}$ as indicated by the simulations. Let us parametrize the temperature-dependent domain wall energy density as follows Kawasaki:2014sqa ; Ferrer:2018uiu
@@
$\rho_{w}(T)=\left\{\begin{matrix}\rho_{w}&\leavevmode\nobreak\ \leavevmode\nobreak\ {T\geq T_{\rm ann}},\\ \rho_{w}\left(\frac{T}{T_{\rm ann}}\right)^{\alpha}&\leavevmode\nobreak\ \leavevmode\nobreak\ {T<T_{\rm ann}},\end{matrix}\right.$ (64)
@@
where $\rho_{w}=\mathcal{A}\sigma_{w}/t$. A fit to the simulation of the domain wall evolution gives $\alpha\simeq 7$ Kawasaki:2014sqa . Then the evolution of the fraction of domain wall energy density is given by
@@
$f_{w}=\left\{\begin{matrix}1.81\times 10^{-3}\frac{\mathcal{A}}{1}\frac{M_{H}(T)}{M_{\odot}}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ \rm GeV}\right)^{3}&M_{H}\leq M_{H}(T_{\rm ann})\\ 1.81\times 10^{-3}\frac{\mathcal{A}}{1}\frac{M_{H}(T)}{M_{\odot}}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ \rm GeV}\right)^{3}\left(\frac{M_{H}(T_{\rm ann})\sqrt{g_{*}(T_{\rm ann})}}{M_{H}(T)\sqrt{g_{*}(T)}}\right)^{\alpha/2}&M_{H}>M_{H}(T_{\rm ann}).\end{matrix}\right.$ (65)
@@
It is worth mentioning that, unlike the PBHs from the collapse of closed domain walls, our mechanism is not sensitive to the evolution of the network after the onset of domain wall annihilation, whose details remain to have a large uncertainty Ferrer:2018uiu .
@@
In Fig. 9, we show the PBH relic abundance $\Omega_{\rm PBH}(f_{w},T)$ from the domain wall number density fluctuations, with the evolution of wall energy fraction according to Eq. (65). The annihilation temperatures $T_{\rm ann}$ and the energy fractions $f_{w}$ at $T_{\rm ann}$ are taken to be (0.1GeV, $(0.1\leavevmode\nobreak\ {\rm GeV},\leavevmode\nobreak\ 0.1)$ (blue curves) and (0.5GeV, $(0.5\leavevmode\nobreak\ {\rm GeV},\leavevmode\nobreak\ 0.2)$ (green curves), which are two benchmarks for PTA observations. The solid and dashed curves correspond to the results with the domain wall number $N=1$ and $N=4$, respectively. The surface tension can be determined by Eq. (55) at $T_{\rm ann}$ and the horizon mass at the annihilation temperature is $M_{H}(T_{\rm ann})=1/(2GH(T_{\rm ann}))$. For our choices, the PBH masses at annihilation are $M(0.1\leavevmode\nobreak\ {\rm GeV})\simeq 0.12M_{\odot}$ (orange vertical line) and $M(0.5\leavevmode\nobreak\ {\rm GeV})\simeq 3.0M_{\odot}$ (red vertical line). We observe that the PBH relic abundance $\Omega_{\rm PBH}$ has a narrow spike around $M(T_{\rm ann})$, below and above which $\Omega_{\rm PBH}$ sharply decreases with the PBH mass. Therefore, PBHs are significantly produced at the moment close to domain wall annihilation. This fact will be adopted to put useful constraints on the model parameter space of the domain wall scenario.
[…]
@@
From Eq. (40) and Fig. 10 we observe that to satisfy various experimental constraints on the PBH relic abundance for $M\sim M_{\odot}$, we should require $\beta(M)\lesssim 10^{-8}-10^{-7}$, corresponding to $x=\delta_{c}/(\sqrt{2}\sigma)\gtrsim 4$. The density fluctuations deep in the tails of the probability distribution exponentially suppress the production of PBHs via the critical density collapse mechanism (see Fig. 10 in appendix D). The parameters that give the probability of a density contrast within $3\sigma$ are strongly constrained by the PBH abundance. To estimate these bounds, we note that the PBH abundance has a sharp spike at the annihilation temperature. Therefore, we calculate the $f_{w}$-dependent critical density contrast $\delta_{c}$ with Eqs. (48) and (49) and determine $\Omega_{\rm PBH}(f_{w}(T_{\rm ann}),T_{\rm ann})$ with Eq. (40) at $T_{\rm ann}$. In table 2, we provide the upper limits on the domain wall energy fraction $f_{w}$ at the annihilation temperature $T_{\rm ann}=0.1$ GeV from the condition $\Omega_{\rm PBH}\leq 10^{-3}$ for various values of the domain wall number $N$. In Fig. 5 and Fig. 6, we show the contours of PBH relic abundance with $\Omega_{\rm PBH}=10^{-3}$ (green solid line) and $\Omega_{\rm PBH}=0.25=\Omega_{\rm DM}$ (gray dot line) in the $f_{w}(T_{\rm ann})-T_{\rm ann}$ plane. We observe from Fig. 5 that the scenario for PTAs with the domain wall number $N=1$ (corresponding to models with a $Z_{2}$ symmetry) has been excluded by the PBH abundance from domain wall number density fluctuations. A marginal space out of $2\sigma$ CL remains available for the case with $N=2$. We therefore also exclude such models for explaining the nano-Hz SGWB. For the case of $N=3$ and $N=4$, most of the parameter space for NG15 and about half of the space region for IPTA-DR2 are excluded by the PBH relic abundance. Finally, for models with $N=6$ and $N=10$, we show in Fig. 6 that most of the parameter space for IPTA-DR2 and about half of the space region for NG15 are not constrained by the PBH formation. Models for $N\sim 10$ domain walls include the clockwork axion, in which several domain walls are produced following the spontaneous breakdown of a set of discrete shift symmetries.
@@ 10 Conclusions and discussions
In this work, we propose that PBHs can be formed from the number density fluctuations of the domain walls in a late stage. The network consisting of a horizon-sized domain wall with a small number density could lead to large Poisson fluctuations. Since the energy density of the domain wall scales as $t^{-1}$ in the scaling regime, the probability for the production of horizon-size density perturbations with large amplitudes from domain wall number density fluctuations grow with time. Once the density contrast exceeds the critical value, it stops expanding and collapses to form PBHs. The outburst of PBH formation takes place when the fraction of domain wall energy density increases to $f_{w}\sim 0.1$. In addition, the domain wall with negative pressure could soften the equation of state and reduce the gas pressure that resists the collapse of density perturbations and, therefore, increase the probability of PBH formation.
By a fit to the NANOGrav and IPTA-DR2 datasets, we show that the gravitational waves from domain wall annihilation could well explain the nano-Hz SGWB if the domain walls account for $\sim 10$% of the total energy density at the annihilation temperature $T_{\rm ann}\sim 0.1$ GeV. Thus, the formation of PBHs from domain wall number density fluctuations before the annihilations can be significant. We find that the overproduction of PBH abundance from the fluctuations of the network with the domain wall numbers $N=1$ and $N=2$ have constrained the parameter space for the interpretation of PTA observations. We therefore exclude the domain wall scenario for the PTAs in models with a $Z_{2}$ symmetry. Some of the parameter space for NG15 is constrained in models with $N\lesssim 10$. With these results, we conclude that scenarios for NG15 in terms of an explicit $Z_{\mathcal{N}}$ symmetry King:2023ayw ; Pham:2024vso in a model with a soft $U(1)$ symmetry breaking term Chiang:2019oms ; Bai:2023cqj
$V(\Phi)\supset\frac{1}{M^{\mathcal{N}-4}}(\Phi^{\mathcal{N}}+\Phi^{\dagger\mathcal{N}}),$ (66)
where $\Phi$ is a complex scalar $\Phi$, $M$ has a mass dimension, and $\mathcal{N}\leq 4$ from the renormalizability, are strongly constrained.
To realize a model with $\mathcal{N}\sim 10$ domains (and the domain wall number $N\sim\mathcal{N}$), one can consider $\mathcal{N}$ copies of complex scalars $\Phi_{i}$ (with $i=1,...,\mathcal{N}$). The global $U(1)^{\mathcal{N}}$ symmetry is spontaneously broken when each of $\Phi_{i}$ obtains a common vacuum expectation value. We can further explicitly break the $\mathcal{N}$ $U(1)$ symmetries to $\mathcal{N}-1$ discrete shift symmetries and one $U(1)$ symmetry by a term $\Phi_{i}^{\dagger}\Phi_{i+1}^{q}$, which then leads to the formation of $\mathcal{N}-1$ domains. Such symmetries have been naturally embedded in the framework of the clockwork axion models Kaplan:2015fuy and deconstruction of extradimensional grand unified theories (GUTs) Arkani-Hamed:2001kyx ; Li:2002xd .
Another realization of the $Z_{\mathcal{N}}$ symmetry is the axion-like models (see, for example, Ref. ZambujalFerreira:2021cte ). Consider a hidden non-Abelian $SU(n)$ gauge sector consisting of a complex scalar $\Phi$ and a set of vector-like quarks. The potential of $\Phi$ possesses a global $U(1)$ symmetry, which is spontaneously broken by the Higgs mechanism. If the vector-like quarks are charged under the $U(1)$ transformation, then the quantum anomaly of $U(1)$ with respect to the $SU(n)$ gauge group can break the $U(1)$ symmetry to its $Z_{\mathcal{N}}$ subgroup via the instanton effect during the confinement of the $SU(n)$ symmetry, which gives rise to an axion-like periodic potential
$V(a)\simeq\Lambda_{n}^{4}\left(1-\cos\left(\frac{\mathcal{N}a}{v}\right)\right),$ (67)
@@
We also note that our constraints can be applied to models with the domain wall number $N\lesssim 10$. For models with the domain wall number $N\gg 10$, the production of PBHs from the domain wall number density fluctuations is strongly suppressed by the Poisson statistics and, therefore, the constraints are insignificant.
@@
Finally, let’s briefly review uncertainties that commonly exist in the critical collapse PBH formation mechanism Carr:2016drx . (i) Non-sphericity: We have assumed a homogeneous and spherically symmetric perturbation. The non-sphericity of the perturbation can improve the threshold and reduce the possibility for PBH formation Kuhnel:2016exn . In our scenario, the non-sphericality of the overdense region induced by the domain wall fluctuations is expected to be small since the radiation always dominates the universe. (ii) Non-Gaussianity: We have assumed a Gaussian probability density distribution for the density contrast. The critical collapse mechanism for PBH formation relies on the exponential suppression in the upper tail of the probability distribution. Therefore, a non-Gaussian component in the distribution could exponentially change the abundance of PBH. Further investigations on the non-Gaussianity in the domain wall fluctuations are needed. Nevertheless, the fluctuations do not affect the CMB-scale observations if the domain walls annihilate long before the onset of BBN.
@@ 11 Acknowledgments
BQL is supported in part by Zhejiang Provincial Natural Science Foundation of China under Grant No. LQ23A050002. CWC is supported in part by the National Science and Technology Council under Grant No. NSTC-111-2112-M-002-018-MY3. TL is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201504, by the Project No. 12275333 supported by the National Natural Science Foundation of China, by the Scientific Instrument Developing Project of the Chinese Academy of Sciences, Grant No. YJKYYQ20190049, and by the International Partnership Program of Chinese Academy of Sciences for Grand Challenges, Grant No. 112311KYSB20210012.
@@ Appendix A Domain wall energy tension in toy model
Let’s consider the following toy model for the real scalar $\phi$ as an example Vilenkin:1984ib ; Gelmini:1988sf ; Saikawa:2017hiv :
$\mathcal{L}=\frac{1}{2}\left(\partial_{\mu}\phi\right)^{2}-V(\phi),$ (68)
@@
$V(\phi)=\frac{1}{4}\lambda\left(\phi^{2}-v^{2}\right)^{2}.$ (69)
The $Z_{2}$ symmetry of the model is spontaneously broken when $\phi$ obtains vacuum expectation values ±$\pm v$. The manifold $\mathcal{M}$ of the degenerate vacua contains two points, $\langle\phi\rangle=+v$ and $\langle\phi\rangle=-v$, which are separated by the domain wall of the false vacuum with $\langle\phi\rangle=0$.
Using the Lagrangian density (68), the description of the domain wall can be obtained by solving the equation of motion
$\square\phi+\lambda\phi\left(\phi^{2}-v^{2}\right)=0,$ (70)
@@
$\phi(z)=v\tanh(x/\delta),$ (71)
where $\delta=1/(\sqrt{\lambda/2}v)$ denotes the thickness of the wall. The energy-momentum tensor for the domain wall solution (71) is given by
$T_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi-g_{\mu\nu}\mathcal{L}.$ (72)
With the wall configuration (71) and $\partial_{\mu}\phi(x)=0$ for $\mu\neq x$, we have
$T_{00}=-T_{22}=-T_{33}=\frac{1}{2}\left(\partial_{x}\phi\right)^{2}+V(\phi)\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ T_{11}=0.$ (73)
@@
$T_{\mu\nu}=\frac{\lambda}{2}v^{4}[{\rm sech}(x/\delta)]^{4}\operatorname{diag}(1,0,-1,-1).$ (74)
@@
[…]
Using Eq. (74), the surface energy density of the wall is given by
$\sigma_{w}=S_{00}=-S_{22}=-S_{33}=\frac{4}{3}\sqrt{\frac{\lambda}{2}}v^{3},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ S_{11}=0.$ (76)

### v2

@@ 10 Constraints and implications from PBH formation
Most of the domain wall energy $\rho_{w}$ is released at the annihilation temperature $T_{\rm ann}$, beyond which $\rho_{w}$ scales as $T^{\alpha}$ as indicated by the simulations. Let us parametrize the temperature-dependent domain wall energy density as follows [77, 41]
@@
$\rho_{w}(T)=\left\{\begin{matrix}\rho_{w}&\leavevmode\nobreak\ \leavevmode\nobreak\ {T\geq T_{\rm ann}},\\ \rho_{w}\left(\frac{T}{T_{\rm ann}}\right)^{\alpha}&\leavevmode\nobreak\ \leavevmode\nobreak\ {T<T_{\rm ann}},\end{matrix}\right$. (10.1)
@@
where $\rho_{w}=\mathcal{A}\sigma_{w}/t$. A fit to the simulation of the domain wall evolution gives $\alpha\simeq 7$ [77]. Then the evolution of the fraction of domain wall energy density is given by
@@
$f_{w}=\left\{\begin{matrix}1.81\times 10^{-3}\frac{\mathcal{A}}{1}\frac{M_{H}(T)}{M_{\odot}}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ \rm GeV}\right)^{3}&M_{H}\leq M_{H}(T_{\rm ann})\\ 1.81\times 10^{-3}\frac{\mathcal{A}}{1}\frac{M_{H}(T)}{M_{\odot}}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ \rm GeV}\right)^{3}\left(\frac{M_{H}(T_{\rm ann})\sqrt{g_{*}(T_{\rm ann})}}{M_{H}(T)\sqrt{g_{*}(T)}}\right)^{\alpha/2}&M_{H}>M_{H}(T_{\rm ann}).\end{matrix}\right$. (10.2)
@@
It is worth mentioning that, unlike the PBHs from the collapse of closed domain walls, our mechanism is not sensitive to the evolution of the network after the onset of domain wall annihilation, whose details remain to have a large uncertainty [41].
@@
In Fig. 9, we show the PBH relic abundance $\Omega_{\rm PBH}(f_{w},T)$ from the domain wall number density fluctuations, with the evolution of wall energy fraction according to Eq. (10.2). The annihilation temperatures $T_{\rm ann}$ and the energy fractions $f_{w}$ at $T_{\rm ann}$ are taken to be $(0.1\leavevmode\nobreak\ {\rm GeV},\leavevmode\nobreak\ 0.1)$ (blue curves) and $(0.5\leavevmode\nobreak\ {\rm GeV},\leavevmode\nobreak\ 0.2)$ (green curves), which are two benchmarks for PTA observations. The solid and dashed curves correspond to the results with the domain wall number $N=1$ and $N=4$, respectively. The surface tension can be determined by Eq. (8.6) at $T_{\rm ann}$ and the horizon mass at the annihilation temperature is $M_{H}(T_{\rm ann})=1/(2GH(T_{\rm ann}))$. For our choices, the PBH masses at annihilation are $M(0.1\leavevmode\nobreak\ {\rm GeV})\simeq 0.12M_{\odot}$ (orange vertical line) and $M(0.5\leavevmode\nobreak\ {\rm GeV})\simeq 3.0M_{\odot}$ (red vertical line). We observe that the PBH relic abundance $\Omega_{\rm PBH}$ has a narrow spike around $M(T_{\rm ann})$, below and above which $\Omega_{\rm PBH}$ sharply decreases with the PBH mass. Therefore, PBHs are significantly produced at the moment close to domain wall annihilation. This fact will be adopted to put useful constraints on the model parameter space of the domain wall scenario.
Figure 9: The PBH relic abundance produced from the number density fluctuations of domain walls with $(T_{\rm ann},\leavevmode\nobreak\ f_{w}(T_{\rm ann}))=(0.1\leavevmode\nobreak\ {\rm GeV},\leavevmode\nobreak\ 0.1)$ (blue curves) and $(0.5\leavevmode\nobreak\ {\rm GeV},\leavevmode\nobreak\ 0.2)$ (green curves). The critical density contrast $\delta_{c}=0.453$. The solid and dashed curves represent the results with $N=1$ and $N=4$, respectively. The red and orange vertical lines represent the PBH mass at the temperature of domain wall annihilation for the two scenarios, respectively.
@@
From Eq. (6.10) and Fig. 11 we observe that to satisfy various experimental constraints on the PBH relic abundance for $M\sim M_{\odot}$, we should require $\beta(M)\lesssim 10^{-8}-10^{-7}$, corresponding to $x=\delta_{c}/(\sqrt{2}\sigma)\gtrsim 4$. The density fluctuations deep in the tails of the probability distribution exponentially suppress the production of PBHs via the critical density collapse mechanism (see Fig. 11 in appendix D). The parameters that give the probability of a density contrast within $3\sigma$ are strongly constrained by the PBH abundance. To estimate these bounds, we note that the PBH abundance has a sharp spike at the annihilation temperature. Therefore, we calculate the $f_{w}$-dependent critical density contrast $\delta_{c}$ with Eqs. (7.1) and (7.2) and determine $\Omega_{\rm PBH}(f_{w}(T_{\rm ann}),T_{\rm ann})$ with Eq. (6.10) at $T_{\rm ann}$. In table 2, we provide the upper limits on the domain wall energy fraction $f_{w}$ at the annihilation temperature $T_{\rm ann}=0.1$ GeV from the condition $\Omega_{\rm PBH}\leq 10^{-3}$ for various values of the domain wall number $N$. In Fig. 5 and Fig. 6, we show the contours of PBH relic abundance with $\Omega_{\rm PBH}=10^{-3}$ (green solid line) and $\Omega_{\rm PBH}=0.25=\Omega_{\rm DM}$ (gray dot line) in the $f_{w}(T_{\rm ann})-T_{\rm ann}$ plane. We observe from Fig. 5 that the scenario for PTAs with the domain wall number $N=1$ (corresponding to models with a $Z_{2}$ symmetry) has been excluded by the PBH abundance from domain wall number density fluctuations. A marginal space out of $2\sigma$ CL remains available for the case with $N=2$. We therefore also exclude such models for explaining the nano-Hz SGWB. For the case of $N=3$ and $N=4$, most of the parameter space for NG15 and about half of the space region for IPTA-DR2 are excluded by the PBH relic abundance. Finally, for models with $N=6$ and $N=10$, we show in Fig. 6 that most of the parameter space for IPTA-DR2 and about half of the space region for NG15 are not constrained by the PBH formation. Models for $N\sim 10$ domain walls include the clockwork axion, in which several domain walls are produced following the spontaneous breakdown of a set of discrete shift symmetries.
Figure 10: The PBH energy fraction for the monochromatic PBH mass spectrum. The colored regions represent the experimental constraints from OGLE [102], LIGO [103], galactic radio observations [104], EDGES experiments [105], CMB spectrum measurements [106], and ultra-faint dwarfs dynamics [107]. The blue, red, and green peaks denote the results for the three benchmarks with $T_{\rm ann}=500$ MeV, 100 MeV, and 70 MeV, respectively. The energy fraction of the domain wall and the domain number are fixed as $f_{w}(T_{\rm ann})=0.125$ and $N=4$.
In Fig. 10, the blue, red, and green peaks represent the PBH energy fraction for the monochromatic PBH mass spectrum, corresponding to three benchmarks that can account for the nano-Hz stochastic gravitational wave background (SGWB) observations. We utilize the PBHbounds package [108] to illustrate various experimental constraints on the PBH energy fraction, as indicated by the colored regions as follows
  • The grey region denotes the most recent OGLE bounds from 20-year observations on PBH-induced microlensing events [102].
  • The yellow area represents constraints from LIGO experiments, including observations on SGWB [109], direct searches for gravitational waves from subsolar masses [109], and direct constraints on PBH-PBH mergers with LIGO [103].
  • The purple region indicates constraints from galactic PBH accretion based on radio observations [104].
  • The light-blue region represents constraints from distortions in the CMB spectrum [106].
  • The green region corresponds to constraints from energy injection in the 21-cm signal by EDGES experiments [105] due to accreting PBHs.
  • The grey region also includes constraints from ultra-faint dwarf dynamics [107].
We observe that the benchmarks denoted by the green and red curves, which align with NANOGrav and IPTA SGWB observations, are well below the PBH abundance constraints. In contrast, the blue curve is marginally constrained by the OGLE bounds.
@@ 11 Conclusions and discussions
In this work, we propose that PBHs can be formed from the number density fluctuations of the domain walls in a late stage. The network consisting of a horizon-sized domain wall with a small number density could lead to large Poisson fluctuations during the horizon crossing of the wall. Once the density contrast exceeds the critical value, it stops expanding and collapses to form PBHs. The PBH formation outburst occurs when the fraction of domain wall energy density increases to $f_{w}\sim 0.1$. In addition, the domain wall with negative pressure could soften the equation of state and reduce the gas pressure that resists the collapse of density perturbations, increasing the probability of PBH formation.
With a fit to the NANOGrav and IPTA-DR2 datasets, we showed that the gravitational waves from domain wall annihilation could well explain the nano-Hz SGWB if the domain walls account for $\sim 10$% of the total energy density at the annihilation temperature $T_{\rm ann}\sim 0.1$ GeV. We also demonstrated that the abundance of PBHs can impose stringent constraints on particle models proposed to explain PTA observations of nano-Hz SGWB. For models with a domain wall number $N\lesssim 2$, the parameter space compatible with PTA observations is excluded due to the overproduction of PBHs. The models with $N\gtrsim 3$ also face strong constraints from the PBH overabundance, significantly limiting their viability.
@@
We note that our constraints are applicable to models with a domain wall number $N\lesssim 10$ (the realizations of such models can be found in appendix F). For models with a domain wall number $N\gg 10$, the production of PBHs from domain wall number density fluctuations is strongly suppressed by Poisson statistics. Consequently, the constraints on such models become insignificant.
@@
Finally, let us briefly review uncertainties that commonly exist in the critical collapse PBH formation mechanism [110]. (i) Non-sphericity: We have assumed a homogeneous and spherically symmetric perturbation. The non-sphericity of the perturbation can improve the threshold and reduce the possibility for PBH formation [111]. In our scenario, the non-sphericality of the overdense region induced by the domain wall fluctuations is expected to be small since the radiation always dominates the Universe. (ii) Non-Gaussianity: We have assumed a Gaussian probability density distribution for the density contrast. The critical collapse mechanism for PBH formation relies on the exponential suppression in the upper tail of the probability distribution. Therefore, a non-Gaussian component in the distribution could exponentially change the abundance of PBH. Further investigations on the non-Gaussianity in the domain wall fluctuations are needed. Nevertheless, the fluctuations do not affect the CMB-scale observations if the domain walls annihilate long before the onset of BBN.
This paper is a joint work of Ref. [112], in which we seek to explore a common primordial origin for both PTA observations and LIGO-observed black hole merger events, taking into account black hole accretion.
@@ 12 Acknowledgments
BQL is supported in part by the National Natural Science Foundation of China under Grant No. 12405058 and by the Zhejiang Provincial Natural Science Foundation of China under Grant No. LQ23A050002. CWC is supported in part by the National Science and Technology Council under Grant No. NSTC-111-2112-M-002-018-MY3. TL is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201504, by the Project No. 12275333 supported by the National Natural Science Foundation of China, by the Scientific Instrument Developing Project of the Chinese Academy of Sciences, Grant No. YJKYYQ20190049, and by the International Partnership Program of Chinese Academy of Sciences for Grand Challenges, Grant No. 112311KYSB20210012.
@@ Appendix A Domain wall energy tension in toy model
Let’s consider the following toy model for the real scalar $\phi$ as an example [45, 46, 113]:
$\mathcal{L}=\frac{1}{2}\left(\partial_{\mu}\phi\right)^{2}-V(\phi),$ (A.1)
@@
$V(\phi)=\frac{1}{4}\lambda\left(\phi^{2}-v^{2}\right)^{2}.$ (A.2)
The $Z_{2}$ symmetry of the model is spontaneously broken when $\phi$ obtains vacuum expectation values $\pm v$. The manifold $\mathcal{M}$ of the degenerate vacua contains two points, $\langle\phi\rangle=+v$ and $\langle\phi\rangle=-v$, which are separated by the domain wall of the false vacuum with $\langle\phi\rangle=0$.
Using the Lagrangian density (A.1), the description of the domain wall can be obtained by solving the equation of motion
$\square\phi+\lambda\phi\left(\phi^{2}-v^{2}\right)=0,$ (A.3)
@@
$\phi(z)=v\tanh(x/\delta),$ (A.4)
where $\delta=1/(\sqrt{\lambda/2}v)$ denotes the thickness of the wall. The energy-momentum tensor for the domain wall solution (A.4) is given by
$T_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi-g_{\mu\nu}\mathcal{L}.$ (A.5)
With the wall configuration (A.4) and $\partial_{\mu}\phi(x)=0$ for $\mu\neq x$, we have
$T_{00}=-T_{22}=-T_{33}=\frac{1}{2}\left(\partial_{x}\phi\right)^{2}+V(\phi)\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ T_{11}=0.$ (A.6)
@@
$T_{\mu\nu}=\frac{\lambda}{2}v^{4}[{\rm sech}(x/\delta)]^{4}\operatorname{diag}(1,0,-1,-1).$ (A.7)
@@
$S_{\mu\nu}=\int dxT_{\mu\nu}.$ (A.8)
Using Eq. (A.7), the surface energy density of the wall is given by
$\sigma_{w}=S_{00}=-S_{22}=-S_{33}=\frac{4}{3}\sqrt{\frac{\lambda}{2}}v^{3},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ S_{11}=0.$ (A.9)